the largest of three consecutive numbers is n. the square of this number exceeds the sum of the other two numbers by 38. find the three numbers.

Answer by mananth(16946)

If n is the largest
then( n-1) & (n-2) are the preceding numbers
n^2=(n-1)+(n-2)+38
n^2=2n-3+38
n^2=2n+35
n^2-2n-35=0
n^2-7n+5n-35=0
n(n-7)+5(n-7)=0
(n-7)(n+5)=0
n= 7 OR -5
7,6,5
OR
-5,-6,-7